# Nearly-polynomial inverse theorem for the $U^{d}$ norm in degree $d+1$

### 17 Mar 2026 v1

@@ 4 $\operatorname{rk}^{*}$
For a polynomial $f$, $\operatorname{rk}^{*}(f)$ is the smallest $r$ such that:
  • $f=\sum_{i=1}^{r}f_{i}$ for some reducible polynomials $f_{i}=\alpha_{i}\beta_{i}$ with $\deg(f_{i})\leq\deg(f)$, and
@@
We have $\operatorname{rk}^{*}(f)=\infty$ if no such decomposition exists, for example if $\deg(f)=1$; however, as $f$ need not be homogeneous, this can happen in higher degree as well (see Remark 4.2 below).
We note that if $f$ is homogeneous then $\operatorname{rk}^{*}(f)=\operatorname{rk}(f)$; indeed, the factors may be assumed to be linearly independent forms (see Lemma 4.6), so the affine independence condition is trivially satisfied.
As an example for $\operatorname{rk}^{*}$, consider a polynomial of the form $f=(A+B)\cdot\ell+B$, where $\deg(A)=\deg(B)>1$ and $\deg(\ell)=1$. While $f$ has the decomposition $f=\ell A+(\ell+1)B$, this is not a $\operatorname{rk}^{*}$-decomposition since its degree-$1$ factors $\ell$ and $\ell+1$ are not affinely independent. In fact, $\operatorname{rk}^{*}(f)$ may be arbitrarily large (see Remark 4.2 next).
@@
Interestingly, $\operatorname{rk}^{*}(f)$ can be more closely related to $\operatorname{bias}(f)$, or rather to analytic rank $\operatorname{AR}(f)=-\log_{p}(\operatorname{bias}(P))$, than other standard notions of rank in the literature. For concreteness, consider the notion used by Green and Tao in [GT09]: $\operatorname{rank}(f)$ is the least number $k$ of polynomials $\alpha_{1},\ldots,\alpha_{k}$ with $\deg(\alpha_{i})<\deg(f)$ such that $f\equiv\Gamma(\alpha_{1},\ldots,\alpha_{k})$ for some function $\Gamma\colon\mathbb{F}^{k}\to\mathbb{F}$. Consider the following polynomials $f$:
  • $f(x,y)=A(x)+\ell(y)$ with $\deg(A)>\deg(\ell)=1$. Clearly, $\operatorname{rank}(f)\leq\operatorname{rank}(A)+1<\infty$. However, $\operatorname{rk}^{*}(f)=\infty$, and moreover, $\operatorname{bias}(f)=\operatorname{bias}(A)\operatorname{bias}(\ell)=0$, so $\operatorname{AR}(f)=\infty$ as well.
  • $f(x,y)=(A-B)\ell+B$ with $\ell=\ell(y)$ of degree $1$, and $A=A(x)$,$B=B(x)$ random of degree $d\geq 2$ on $n$ variables. Clearly, $\operatorname{rank}(f)\leq 3$. In contrast, we claim that both $\operatorname{rk}^{*}(f)$ and $\operatorname{AR}(f)$ grow with $n$.
    Consider a $\operatorname{rk}^{*}$-decomposition of $f$. By affine independence, setting all degree-$1$ factors to 0 gives a non-empty affine subspace. Let $U$ be the non-empty intersection of this affine subspace with the hypersurface $\ell=c$, for an appropriate $c\in\mathbb{F}$. Then $\operatorname{codim}(U)\leq\operatorname{rk}^{*}(f)+1$, and the restriction $f^{\prime}$ of $f$ to $U$ has $\operatorname{rank}(f^{\prime})\leq\operatorname{rk}^{*}(f)$. By construction, $f^{\prime}$ is a nonzero linear combination of $A$ and $B$ restricted to $U$, so $\operatorname{rank}(f^{\prime})$ grows with $n$, and thus also $\operatorname{rk}^{*}(f)$.
    As for $\operatorname{AR}(f)$, we have $\operatorname{bias}(f)=\operatorname*{\mathbb{E}}_{x}\chi(B(x))\operatorname*{\mathbb{E}}_{y}\chi(\ell(y)(A(x)-B(x)))=\operatorname*{\mathbb{E}}_{x}\chi(B(x))\mathbf{1}_{A=B}(x)\\ =\sum_{b}\chi(b)\mathbb{P}(B=b,A=b)$, so $\operatorname{bias}(f)\approx(1/q)\sum_{b}\chi(b)\mathbb{P}(B=b)=(1/q)\operatorname{bias}(B)$, implying that $\operatorname{AR}(f)\approx n$, as claimed.
@@ 4.1 Properties of $\operatorname{rk}^{*}$
$\operatorname{rk}^{*}(g+h)\leq\operatorname{rk}^{*}(g)+\operatorname{rk}^{*}(h)$ if $\deg(h)\neq\deg(g)$.
@@
That being said, we show that, perhaps surprisingly, adding to a polynomial a partial derivative does not increase its $\operatorname{rk}^{*}$.
@@
For any form $g$ and any direction vector $c$,
$\operatorname{rk}^{*}(g+\partial_{c}g)\leq 2\operatorname{rk}(g).$
@@
For any polynomial $g+h$ with top-degree component $g$, and any direction vector $c$,
$\operatorname{rk}^{*}(g+h)\leq 2\operatorname{rk}(g)+\operatorname{rk}^{*}(h-\partial_{c}g).$
We now prove the above properties of $\operatorname{rk}^{*}$.
@@
Put $r_{1}=\operatorname{rk}^{*}(g)$, $r_{2}=\operatorname{rk}^{*}(h)$. Let $g=\sum_{i=1}^{r_{1}}g_{i}$ and $h=\sum_{i=1}^{r_{2}}h_{i}$, with $g_{i}$ and $h_{i}$ reducible polynomials with $\deg(g_{i})\leq\deg(g)$ and $\deg(h_{i})\leq\deg(h)$, be decompositions witnessing $\operatorname{rk}^{*}(g)$ and $\operatorname{rk}^{*}(h)$, respectively. We have $g+h=\sum_{i=1}^{r_{1}}g_{i}+\sum_{i=1}^{r_{2}}h_{i}$. Assume without loss of generality that $\deg(h)<\deg(g)$. Since $\deg(h_{i})\leq\deg(h)$, observe that the affine independence condition for $\operatorname{rk}^{*}(g+h)$ applies only for the $g_{i}$, and is therefore satisfied by the definition of $\operatorname{rk}^{*}(g)$. Thus, $\operatorname{rk}^{*}(g+h)\leq r_{1}+r_{2}=\operatorname{rk}^{*}(g)+\operatorname{rk}^{*}(h)$, as needed. ∎
@@
Put $r=\operatorname{rk}(g)$, and let $g=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ be a $\operatorname{rk}$-decomposition. Suppose for contradiction that $\alpha_{r}=\sum_{i=1}^{r-1}c_{i}\alpha_{i}$ with $c_{i}$ scalars. By homogeneity, we may assume that $\deg(\alpha_{i})=\deg(\alpha_{r})$ whenever $c_{i}\neq 0$. Observe that $g$ has the decomposition $g=\sum_{i=1}^{r-1}\alpha_{i}(\beta_{i}+c_{i}\beta_{r})$. Each summand is a reducible form, since $\alpha_{i}$ is homogeneous as well as $\beta_{i}+c_{i}\beta_{r}$; indeed, $\deg(\beta_{i})=k-\deg(\alpha_{i})=k-\deg(\alpha_{r})=\deg(\beta_{r})$ whenever $c_{i}\neq 0$. Thus, $r=\operatorname{rk}(g)\leq r-1$, a contradiction. ∎
@@
Put $r=\operatorname{rk}(g)$ and $k=\deg(g)$, and write $g=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ with forms $\alpha_{i},\beta_{i}$ satisfying $\deg(\alpha_{i})\leq\deg(\beta_{i})<k$ and $\deg(\alpha_{i}\beta_{i})=k$.
@@
Now, since $\deg(\alpha_{i}+\partial_{c}\alpha_{i})=\deg(\alpha_{i})\leq\deg(\beta_{i})=\deg(\beta_{i}+\partial_{c}\beta_{i})$, the linear components of the factors $\alpha_{i}+\partial_{c}\alpha_{i}$ are precisely the linear factors $\alpha_{i}$ in the above decomposition for $g$, which by Lemma 4.6 are linearly independent. As for the remaining terms, note that $\deg((\partial_{c}\alpha_{i})(\partial_{c}\beta_{i}))\leq\deg(\alpha_{i}\beta_{i})-2=k-2<\deg(g+\partial_{c}g)$, so $\operatorname{rk}^{*}(g+\partial_{c}g)$ has no requirement on their (linear) factors. We deduce that the above (non-homogeneous) decomposition is a $\operatorname{rk}^{*}$-decomposition, so $\operatorname{rk}^{*}(g+\partial_{c}g)\leq 2r=2\operatorname{rk}(g)$, as needed.
@@
We have
$\operatorname{rk}^{*}(g+h)$ $=\operatorname{rk}^{*}\big((g+\partial_{c}g)+(h-\partial_{c}g)\big)$
$\leq\operatorname{rk}^{*}(g+\partial_{c}g)+\operatorname{rk}^{*}(h-\partial_{c}g)\leq 2\operatorname{rk}(g)+\operatorname{rk}^{*}(h-\partial_{c}g)$
where the first inequality uses the subadditivity of $\operatorname{rk}^{*}$ in Observation 4.3, and the second inequality uses Lemma 4.4. ∎

### 30 Apr 2026 v2

@@ 4 $\operatorname{rk}^{*}$
For a non-constant polynomial $f$, $\operatorname{rk}^{*}(f)$ is the smallest $r$ such that:
  • $f=\sum_{i=1}^{r}f_{i}$ for some reducible polynomials $f_{i}=\alpha_{i}\beta_{i}$ with $\deg(f_{i})\leq\deg(f)$, [^17] That is, $\deg(\alpha_{i}),\deg(\beta_{i})\geq 1$ and $\deg(\alpha_{i})+\deg(\beta_{i})\leq\deg(f)$. Note that $\deg(\alpha_{i}),\deg(\beta_{i})<\deg(f)$. and
@@
If $f$ is a constant polynomial, we set $\operatorname{rk}^{*}(f)=0$. By convention, $\operatorname{rk}^{*}(f)=\infty$ if no such decomposition exists. This happens when $\deg(f)=1$, and can also happen when $\deg(f)=2$; for example, for any irreducible univariate quadratic. [^18] Affine independence implies at most $n$ top-degree summands having a degree-$1$ factor. For a univariate quadratic, this means at most $n=1$ summands. On the other hand, $\deg(f)\geq 3$ implies $\operatorname{rk}^{*}(f)<\infty$; in fact, $\operatorname{rk}^{*}(f)\leq n+1$. [^19] Indeed, let $q$ be any reducible quadratic with affine part $q_{\leq 1}=f_{\leq 1}$ (say $q(x)=(x_{1}+1)(f_{\leq 1}(x)-f_{0}x_{1})$). Then fqf-q has only monomials of degree at least $2$, so it has a $\operatorname{rk}^{*}$-decomposition with $\alpha_{1}=x_{1},\ldots,\alpha_{n}=x_{n}$. Since $\deg(q)<3\leq\deg(f)$, this gives a $\operatorname{rk}^{*}$-decomposition for $f$ of length at most $n+1$. (This bound cannot be improved to $n$ in general; e.g., any irreducible univariate $f$, assuming $\mathbb{F}$ is not algebraically closed, has $\operatorname{rk}^{*}(f)\geq 2=n+1$.)
Note that if $f$ is homogeneous then $\operatorname{rk}^{*}(f)=\operatorname{rk}(f)$. Indeed, $\operatorname{rk}^{*}(f)\leq\operatorname{rk}(f)$ since the factors may be assumed to be linearly independent forms (see Lemma 4.6), so the affine independence condition trivially holds; conversely, $\operatorname{rk}(f)\leq\operatorname{rk}^{*}(f)$ since for each summand $\alpha_{i}\beta_{i}$, the lower-degree components of $\alpha_{i}$ and of $\beta_{i}$ contribute only monomials of degree below $\deg(f)$ to the product, and thus may all be discarded. Note too that $f$ of positive degree is irreducible if and only if $\operatorname{rk}^{*}(f)>1$, similarly to $\operatorname{rk}$.
The importance of $\operatorname{rk}^{*}(f)$ stems from the fact that it is closely related to correlation with polynomials of degree $\deg(f)-2$, whereas the usual $\operatorname{rk}(f)$ only gives information about correlation with degree $\deg(f)-1$.
As an example for $\operatorname{rk}^{*}$, consider a polynomial of the form $f=(A-B)\cdot\ell+B$, where $\deg(A)=\deg(B)>1$ and $\deg(\ell)=1$. While $f$ has the decomposition $f=\ell A+(-\ell+1)B$, this is not a $\operatorname{rk}^{*}$-decomposition since its degree-$1$ factors $\ell$ and $\ell+1$ are not affinely independent. In fact, $\operatorname{rk}^{*}(f)$ may be arbitrarily large. Thus, it is not surprising that $f$ may have negligible correlation with polynomials of degree at most $\deg(f)-2$ (see Remark 4.2 next). This is in sharp contrast to polynomials with low $\operatorname{rk}^{*}$, which do have significant correlation with polynomials of degree at most $\deg(f)-2$ (see Proposition 4.7 below).
@@
Interestingly, $\operatorname{rk}^{*}(f)$ can be more closely related to $\operatorname{bias}(f)$, or rather to analytic rank $\operatorname{AR}(f)=-\log_{|\mathbb{F}|}(\operatorname{bias}(f))$, than other standard notions of rank in the literature. For concreteness, consider the notion used by Green and Tao in [GT09]: $\operatorname{rank}(f)$ is the least number $k$ of polynomials $\alpha_{1},\ldots,\alpha_{k}$ with $\deg(\alpha_{i})<\deg(f)$ such that $f\equiv\Gamma(\alpha_{1},\ldots,\alpha_{k})$ for some function $\Gamma\colon\mathbb{F}^{k}\to\mathbb{F}$. Consider the polynomial $f(x,y)=(A-B)\ell+B$ with $\ell=\ell(y)$ of degree $1$, and $A=A(x)$, $B=B(x)$ random of constant degree $d\geq 2$ on $n$ variables. Clearly, $\operatorname{rank}(f)\leq 3$. In contrast, we will next show (slightly informally) that $\operatorname{rk}^{*}(f)$ and $\operatorname{AR}(f)$ grow with $n$, the number of $x$ variables. We note that a random $n$-variate polynomial over $\mathbb{F}_{q}$ of degree $d$ has exponential small $q^{-\Omega_{d}(n)}$ correlation with lower-degree polynomials [BHL12, BGY20].
Consider a $\operatorname{rk}^{*}$-decomposition $f=\sum_{i=1}^{\operatorname{rk}^{*}(f)}\alpha_{i}\beta_{i}$. By affine independence, setting to 0 all degree-$1$ factors $\alpha_{i}$ coming from top-degree summands gives an affine subspace that is not empty. Let $U$ be the non-empty intersection of this affine subspace with the affine hyperplane $\ell=c$, for an appropriate $c\in\mathbb{F}$. Note that $\operatorname{codim}(U)\leq\operatorname{rk}^{*}(f)+1$. Observe that the restriction $f^{\prime}$ of $f$ to $U$ has (in the sense of Green-Tao) $\operatorname{rank}(f^{\prime})\leq 2\operatorname{rk}^{*}(f)$, since $\deg(f^{\prime})=d$ (with high probability) and it inherits a decomposition in terms of polynomials of degree at most $\deg(f)-2=d-1$. By construction, $f^{\prime}$ is a nonzero linear combination of $A$ and $B$ restricted to $U$; using our bound on $\operatorname{codim}(U)$ and the randomness of $A$ and $B$, $\operatorname{rank}(f^{\prime})$ grows with $n$, and thus also $\operatorname{rk}^{*}(f)$.
We will next see that $\operatorname{cor}(f,P)\leq|\mathbb{F}|^{-\Omega(n)}$ for any $P\in\operatorname{Poly}_{\deg(f)-2}$. Note that
$\operatorname{cor}(f,P)=|\operatorname*{\mathbb{E}}_{y}\operatorname*{\mathbb{E}}_{x}\chi(f(x,y)-P(x,y))|\leq\max_{y}|\operatorname*{\mathbb{E}}_{x}\chi(f(x,y)-P(x,y))|=\max_{y}\operatorname{cor}(f_{y},P_{y})$
where $f_{y}(x)=f(x,y)$ and $P_{y}(x)=P(x,y)$. For every fixed $y$, we have by construction that $f_{y}$ is a nontrivial linear combination of two random polynomials of degree $d$, and thus is a random polynomial of degree $d$ as well; therefore, since $\deg(P_{y})\leq\deg(P)\leq d-1$, we have that $\operatorname{cor}(f_{y},P_{y})\leq|\mathbb{F}|^{-\Omega(n)}$. By the above, the same follows for $\operatorname{cor}(f,P)$. [^20] An alternative proof that $\operatorname{rk}^{*}(f)$ grows with $n$ follows immediately by combining this correlation upper bound together with Proposition 4.7 below.
Finally, note that this correlation upper bound in particular applies to $\operatorname{bias}(f)=\operatorname{cor}(f,0)$, so $\operatorname{AR}(f)\geq\Omega(n)$, similarly to $\operatorname{rk}^{*}(f)$.
@@ 4.1 Properties of $\operatorname{rk}^{*}$
$\operatorname{rk}^{*}(g+h)\leq\operatorname{rk}^{*}(g)+\operatorname{rk}^{*}(h)$ for positive $\deg(h)\neq\deg(g)$.
@@
That being said, we show that, perhaps surprisingly, adding to a polynomial a partial derivative does not increase its $\operatorname{rk}^{*}$, up to a polynomial of lower degree. [^21] Alternatively, one could omit the lower degree polynomial $\gamma$ from Lemma 4.4, and from Proposition 4.5 below, by redefining $\operatorname{rk}^{*}(f)$ to ignore the components of $f$ of degree at most $\deg(f)-2$.
@@
For any $g\in\operatorname{Form}_{k}$ with $k\geq 2$, and any direction vector $c$,
$\operatorname{rk}^{*}(g+\partial_{c}g+\gamma)\leq\operatorname{rk}(g)$
for some $\gamma\in\operatorname{Poly}_{k-2}$.
For example, for the univariate $g(x)=x^{2}\in\operatorname{Form}_{2}$, we have $\partial_{c}g=2cx$, and it is easy to see that there is a constant $\gamma\in\mathbb{F}$ such that $g+\partial_{c}g+\gamma$ is reducible (is of $\operatorname{rk}^{*}$ 11); indeed, take $\gamma=c^{2}$. Lemma 4.4 can be thought of as a generalization of such identities.
@@
For any $g\in\operatorname{Form}_{k}$ with $k\geq 2$ and $h\in\operatorname{Poly}_{<k}$, and any direction vector $c$,
$\operatorname{rk}^{*}(g+h+\gamma)\leq\operatorname{rk}(g)+\operatorname{rk}^{*}(h-\partial_{c}g)$
for some $\gamma\in\operatorname{Poly}_{k-2}$.
We now prove the above properties of $\operatorname{rk}^{*}$. We start with subadditivity.
@@
Put $r_{1}=\operatorname{rk}^{*}(g)$, $r_{2}=\operatorname{rk}^{*}(h)$, and assume both are finite as otherwise there is nothing to prove. Let $g=\sum_{i=1}^{r_{1}}g_{i}$ and $h=\sum_{i=1}^{r_{2}}h_{i}$, with $g_{i}$ and $h_{i}$ reducible polynomials with $\deg(g_{i})\leq\deg(g)$ and $\deg(h_{i})\leq\deg(h)$, be decompositions witnessing $\operatorname{rk}^{*}(g)$ and $\operatorname{rk}^{*}(h)$, respectively (recall that $g$ and $h$ are not constant). We have $g+h=\sum_{i=1}^{r_{1}}g_{i}+\sum_{i=1}^{r_{2}}h_{i}$. Assume without loss of generality that $\deg(h)<\deg(g)$. Since $\deg(h_{i})\leq\deg(h)$, observe that the affine independence condition for $\operatorname{rk}^{*}(g+h)$ applies only for the $g_{i}$ (of degree $\deg(g)$ with a degree-$1$ factor), and is therefore satisfied by the definition of $\operatorname{rk}^{*}(g)$. Thus, $\operatorname{rk}^{*}(g+h)\leq r_{1}+r_{2}=\operatorname{rk}^{*}(g)+\operatorname{rk}^{*}(h)$, as needed. ∎
We will use the following easy observation.
@@
Put $r=\operatorname{rk}(g)$, and let $g=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ be a $\operatorname{rk}$-decomposition. Suppose for contradiction that $\alpha_{r}=\sum_{i=1}^{r-1}c_{i}\alpha_{i}$ with $c_{i}$ scalars. By homogeneity, we may assume that $\deg(\alpha_{i})=\deg(\alpha_{r})$ whenever $c_{i}\neq 0$. Observe that $g$ has the decomposition $g=\sum_{i=1}^{r-1}\alpha_{i}(\beta_{i}+c_{i}\beta_{r})$. Each summand is a reducible form, since $\alpha_{i}$ is homogeneous as well as $\beta_{i}+c_{i}\beta_{r}$; indeed, $\deg(\beta_{i})=\deg(g)-\deg(\alpha_{i})=\deg(g)-\deg(\alpha_{r})=\deg(\beta_{r})$ whenever $c_{i}\neq 0$. Thus, $r=\operatorname{rk}(g)\leq r-1$, a contradiction. ∎
@@
Put $r=\operatorname{rk}(g)$ and write $g=\sum_{i=1}^{r}\alpha_{i}\beta_{i}$ with forms $\alpha_{i},\beta_{i}$ satisfying $1\leq\deg(\alpha_{i})\leq\deg(\beta_{i})<k$ and $\deg(\alpha_{i}\beta_{i})=k$.
@@
Put $\alpha_{i}^{\prime}=\alpha_{i}+\partial_{c}\alpha_{i}$ and $\beta_{i}^{\prime}=\beta_{i}+\partial_{c}\beta_{i}$, and note that $\deg(\alpha_{i}^{\prime})=\deg(\alpha_{i})\leq\deg(\beta_{i})=\deg(\beta_{i}^{\prime})$. Thus, the linear components of those factors $\alpha^{\prime}_{i}$ with $\deg(\alpha^{\prime}_{i})=1$ are precisely the $\alpha_{i}$ with $\deg(\alpha_{i})=1$, which by Lemma 4.6 are linearly independent. As for the remaining terms, note that $\deg((\partial_{c}\alpha_{i})(\partial_{c}\beta_{i}))\leq\deg(\alpha_{i}\beta_{i})-2=k-2$, so their sum $\gamma$ has $\deg(\gamma)\leq k-2$. Thus, we have a (non-homogeneous) $\operatorname{rk}^{*}$-decomposition of $g+\partial_{c}g+\gamma$, and thus $\operatorname{rk}^{*}(g+\partial_{c}g+\gamma)\leq r=\operatorname{rk}(g)$, as needed. ∎
@@
Apply Lemma 4.4 on $g\in\operatorname{Form}_{k}$ and $c$ to obtain $\gamma\in\operatorname{Poly}_{k-2}$ with $\operatorname{rk}^{*}(g+\partial_{c}g+\gamma)\leq\operatorname{rk}(g)$. If $h-\partial_{c}g$ is a constant $a$, then we are done since $\gamma-a\in\operatorname{Poly}_{k-2}$ and
$\operatorname{rk}^{*}(g+h+(\gamma-a))=\operatorname{rk}^{*}(g+\partial_{c}g+\gamma)\leq\operatorname{rk}(g)=\operatorname{rk}(g)+\operatorname{rk}^{*}(h-\partial_{c}g).$
Otherwise,
$\operatorname{rk}^{*}(g+h+\gamma)$ $=\operatorname{rk}^{*}\big((g+\partial_{c}g+\gamma)+(h-\partial_{c}g)\big)$
$\leq\operatorname{rk}^{*}(g+\partial_{c}g+\gamma)+\operatorname{rk}^{*}(h-\partial_{c}g)\leq\operatorname{rk}(g)+\operatorname{rk}^{*}(h-\partial_{c}g)$
where the first inequality uses the subadditivity of $\operatorname{rk}^{*}$ in Observation 4.3, as the two polynomials have different positive degrees: $k$ and at most k1k-1. ∎